OpenStudy (anonymous):

what does it mean to solve the equation for M then find the value of M in each N

OpenStudy (anonymous):

do you have an example to go by?

OpenStudy (anonymous):

yes hold on

OpenStudy (anonymous):

\[M+3n=7n\]

OpenStudy (anonymous):

\[-2,0,1\]

OpenStudy (anonymous):

that is the end of it

OpenStudy (phi):

solve for M means "get M alone on one side" of the equal sign

OpenStudy (anonymous):

i know that part i ment the part about the N

OpenStudy (phi):

what do you get when you solve for M in this: M+3n=7n

OpenStudy (anonymous):

ok so to solve for M in M+3n=7n you have to leave M by itself M=7n-3n then once you have M by itself you can plug in what you got for M then solve 7n-3n+3n=7n everything cancels out therefore it means that it is exact opposite

OpenStudy (anonymous):

do you understand what i tried to say? if not let me know

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

when u solve for m you have 7n-3n right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then it ends up being M=10 right

OpenStudy (anonymous):

ok now plug in what you got for M you end up with 7n-3n+3n=7n

OpenStudy (anonymous):

how did u get 10?

OpenStudy (anonymous):

i divdied 10 from each side

OpenStudy (anonymous):

never mind that the wrong prblem

OpenStudy (anonymous):

ugh i suck at this

OpenStudy (anonymous):

@mathway

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

Is it n= -2, 0, 1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\(M+3n=7n\) Subtract 3n on both sides. What would you get?

OpenStudy (anonymous):

with the M or with the 3 and 7

OpenStudy (anonymous):